6x2-5x-21=0 solve by the method of factorization.
Answer by Simnepi(216)

factors of -21 are +-3 and +-7
factors of 6 are 2 and 3
by trial and error you find that the coefficient of x
-5 = -14 + 9 (so 2*-7 + 3*3 = -5)
so (2x+3)(3x-7)=0
therefore x = -3/2 or x = 7/3